In the laboratory a "coffee cup" calorimeter, or constant pressure calorimeter, is frequently used to determine the specific heat of a solid, or to measure the energy of a solution phase reaction.

A student heats 64.65 grams of tungsten to 99.08 °C and then drops it into a cup containing 82.26grams of water at 24.43 °C. She measures the final temperature to be 26.31 °C.

The heat capacity of the calorimeter (sometimes referred to as the calorimeter constant) was determined in a separate experiment to be 1.87 J/°C.

Assuming that no heat is lost to the surroundings calculate the specific heat of tungsten.
Answer #1

Heat lost by tungsten= heat gained by water+ heat gained by calorimeter

MCdt= mcdt'+Hdt'

Where

M= mass of tungsten= 64.65 g

C= specific heat capacity of tungsten= ?

dt= change in temperature of tungsten= 99.08-26.31= 72.77 oC

m= mass of water= 82.26 g

c= specific heat of water= 4.186 J/(goC)

dt'= change in temperature of water= 26.31-24.43= 1.88 oC

H= heat capacity of calorimeter= 1.87J/oC

Plug the values we get C = 0.138 J/goC